OpenStudy (anonymous):

simplify the expression. (m^4)(m^6)

OpenStudy (anonymous):

(m^4)(m^6)=m^10

OpenStudy (anonymous):

thanks! i knew something was wrong with my answer.

OpenStudy (kropot72):

\[m ^{4}\times m^{6} =m ^{(4+6)}=m ^{10}\]

OpenStudy (anonymous):

You're welcome!